what is the weight of a block of ice 24in by 24in by 24in, if ice weighs 92 percent as much as water, and water weighs 62.5pounds per cubic feet?kindly show the solutions please.......thank you.
Answer by mananth(16946)

24 in = 2ft
..
volume of ice = 2*2*2 = 8 cu.ft
1cu.ft = 62.5 pounds of water
ice is= 92 % by weight of water
..
weight = 8 * 62.5 * 0.92
=460 pounds
...
m.ananth@hotmail.ca